# Field-based Molecule Generation

### v1

@@ 5.1 QM9
Model Atom Mol Validity Nov (↑)
EDM($\dagger$) 98.1 81.9 91.6 65.7
GDM-AUG 97.6 71.6 90.4 74.6
EDM-Bridge 98.8 84.6 92.0 65.7
MolDiff(*, $\ddagger$) - - 97.0 -
MiDi($\dagger$) 98.3 84.5 96.7 54.1
G-Schnet 95.7 68.1 85.5 -
GeoLDM 98.9 89.4 93.8 -
VoxMol(*) 99.2 89.3 98.7 -
FMG 99.4$\pm$ 0.1 91.5$\pm$ 1.4 98.8 $\pm$ 0.3 68.8 $\pm$ 0.3
QM9 99.9 99.6 99.6 -
Table 2: Molecule graph quality of generated samples from models trained on the QM9 dataset. We report mean and standard deviation over three runs, with 10k samples per model.
Model MST (↑) TVa(↓) TVb(↓)
EDM($\dagger$) 0.57 0.49 0.98
MiDi($\dagger$) 0.55 0.19 0.81
FMG 0.54$\pm$ 0.00 0.52 $\pm$ 0.02 0.95 $\pm$ 0.01
5.2 QM9 conditional generation
We employ the same setting as in previous work (Guan et al., 2023), where we split molecular properties in one thousand equal bins, and form, based on training data statistics, the conditional and marginal categorical distributions $p(N_{a}|\mathbf{c}_{i})$, and $p(\mathbf{c}_{i})$, where $\mathbf{c}_{i}$ is the $i^{\text{th}}$ property bin and $N_{a}$ is the number of atoms. We test our conditional model by sampling two thousand pairs of $[\mathbf{c}_{i},N_{a}]\sim p(\mathbf{c}_{i})p(N_{a}|\mathbf{c}_{i})$, and further sample the actual conditional property $c\sim\mathcal{U}(\mathbf{c}_{i,0},\mathbf{c}_{i,1})$, with which we condition the generative model on $\boldsymbol{\epsilon}_{\theta}(\mathbf{f}_{t},t,N_{a},c)$.
The results and further details of the experiment are shown in Appendix A.2, where we compare the difference between the highest and lowest occupied molecular energies $\Delta\epsilon$ and the polarizability $\alpha$ of the generated molecules, compared to the true, conditional values. FMG outperforms EDM in terms of molecular stability, while EDM is better in its fidelity to the conditioning variables $c$.
Additionally, our method generates similar fidelity to the conditioning variable $c$, even without classification-free guidance on $N_{a}\sim p(N_{a}|c)$. This could prove useful for future situations where the conditional $p(N_{a}|c)$, cannot be accurately determined based on data statistics (e.g., when conditioning on multiple properties $c_{1},c_{2}$, where data may be insufficient to determine $p(N_{a}|c_{1},c_{2})$ accurately).
5.3 GEOM
The Geometric Ensemble Of Molecules (GEOM) dataset (Axelrod & Gómez-Bombarelli, 2022) contains over 450 thousand molecules of up to 181 atoms and 37 million conformations along with their corresponding energies. Following previous work (Guan et al., 2023; Peng et al., 2023a), we select the 30 lowest energy conformers for each molecule graph. Similar to (Peng et al., 2023a), we also restrict the data points to only those molecules that contain atoms from {H, C, N, O, F, P, S, Cl}. Finally, we extract those molecules that can be contained in the pre-defined evaluated positions that determine the tensors $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$, similar to our QM9 experiments. After filtering, the data contains 96.7% of the training molecular conformations and 99.7% unique molecular graph coverage.
[…]
Finally, we analyze the conformational properties of our method in Appendix A.1. Even on this significantly more complex dataset, our method generates most molecular geometries very well.
Table 3: Basic molecule properties for molecular graphs generated by models trained on the GEOM dataset. We report the mean result of 10k generated samples, for both explicit and implicit H models. An H checkmark means the method explicitly generates the H atoms. All properties are expressed in percentages.
Neutrality (↑)
Model H Atom Molecule Val Nov(↑)
EDM($\dagger$) 81.0 0.5 3.0 100
MiDi($\dagger$,$\ddagger$ , ‡) 96.8 31.3 62.7 100
GeoLDM 84.4 - 99.3 -
VoxMol(*) 98.1 75.0 93.4 -
FMG 94.7 16.5 64.9 100
GEOM 99.4 86.2 89.6 -
MolDiff($\dagger$, $\ddagger$) - - 85.0 100
FMG - - 80.8 100
Table 4: Molecule graph quality of generated samples from models trained on the GEOM dataset. We report the mean result of 10k generated samples, for both explicit and implicit H models.
Model H MST(↑) TVa(↓) TVb(↓)
EDM($\dagger$) 0.46 1.51 5.72
MiDi($\dagger$,$\ddagger$) 0.53 0.76 -
FMG 0.49 2.3 4.78
MolDiff($\dagger$, $\ddagger$) 0.62 2.03 -
FMG 0.50 0.64 4.58
5.4 Chirality analysis
We aim to also empirically attest that our proposed method can capture all molecule chiral properties, while current implementations of SO(3) invariant NN lead to invariance w.r.t. enantiomers. We create a simplified experiment, where we attempt to overfit on 48 unique molecules having a carbon tetrahedral configuration connected to O, N, H, and C. We generated the molecule atom positions using RDKit, and analyzed the distribution of the determinants of the atom positions $\mathbf{m}$ of O, N, and H.
We consider matrices defined by $R=[\mathbf{m}_{\texttt{O}}-\mathbf{m}_{\texttt{C}},\mathbf{m}_{\texttt{N}}-\mathbf{m}_{\texttt{C}},\mathbf{m}_{\texttt{H}}-\mathbf{m}_{\texttt{C}}]^{T}\in\mathbb{R}^{3\times 3}$, where $\mathbf{m}_{C}$ is the central carbon atom of the tetrahedral configuration, and the triplets $(\mathbf{m}_{O},\mathbf{m}_{C},\mathbf{m}_{H})$ are the positions of three of the four atoms it connects to. The enantiomer configuration of the generated and the data molecules can be uniquely identified by the sign of $\operatorname{det}R$, while the volume between the four atoms is defined by its absolute value $|\operatorname{det}R|$. In Figure 5, the determinant values of molecules generated by our method match those found in the data, while the equivariant method (EDM) has any of the two enantiomers equally likely. We further assess the SO(3) variant method defined by (Guan et al., 2023), which concatenates atom positions to its GNN node states, that should, in theory, be able to capture enantiomer states. Surprisingly, although the method now clearly starts recognizing the two different enantiomers, it fails to produce realistic conformations and generates many molecules with approximately 0 volume, i.e., the central tetrahedral carbon C and the three connecting atoms, N, H, and O, are approximately coplanar.
Please refer to Appendix A.3 for details of this experiment and theoretical analysis of the current SO(3) implementations’ behavior related to enantiomer configurations.
Figure 5: Kernel density estimation of the triplet vector $[\mathbf{m}_{\texttt{O}}$ $\mathbf{m}_{\texttt{C}}$, $\mathbf{m}_{\texttt{N}}$ $\mathbf{m}_{\texttt{C}}$, $\mathbf{m}_{\texttt{H}}$-$\mathbf{m}_{\texttt{C}}]^{T}$ determinants, resulted from the data and generated molecules. EDM is unable to distinguish and correctly generate the enantiomer distribution, while the SO(3) variant method (GNN) generates improper, 0 volume conformations. FMG matches the enantiomer distribution.
5.5 Ablation
[…]
For FMGdist, we use the same lookup table from (Guan et al., 2023) to determine bonds based on atom distances. In Table 5, the performance of FMGdist is remarkably good. This further confirms that our method generates very good conformations (having bond distances per type within acceptable margins) since it would otherwise be impossible for this method to have such high atom and molecule stability.
Table 5: FMG performance with various design choices. All basic molecular properties (Val and Neutrality Atom and Mol) are expressed in percentages.
Neutrality
Model Val Atom Mol TVa(↓) TVb(↓)
@@ 6 Conclusion
In this work, we show how continuous field-based molecular representations can be successfully employed for drug-like molecule generation. We achieve highly competitive molecule stability and analyze enantiomers, a previously omitted molecular property. We show how our method can capture this molecular property, while point-cloud methods cannot, and fail even in the SO(3) variant context.
@@ 7 Acknowledgements
We acknowledge the computational resources provided by the Aalto Science-IT Project.
@@ A.1 Conformational analysis
We report the Wasserstein distances between the generated and data bond angle distributions in Table 6. We calculated the Wasserstein W1 bond angle (BA) and bond length (BL) distances for each atom and bond type and respectively averaged over atom and bond types (please refer to Appndix H for details). In Figure 6, we also plot the CDF of the distributions, based on which the W1 distances were computed.
Note that for the GEOM dataset, MiDi uses explicitly generated aromatic bond types, and thus has a different bond type set $\mathcal{B}$ compared to ours. This also affects the distributions of the bond types 1 and 2 and we therefore omit MiDi from the GEOM bond distance metric BL.
[…]
GEOM QM9
[…]
EDM($\dagger$) 9.09 5.48 0.47 0.68
MiDi($\dagger$,$\ddagger$ , ‡) - 1.55 1.96 0.95
FMG 2.99 1.99 0.52 $\pm$ 0.02 0.95 $\pm$ 0.01
Figure 6: Cumulative distribution function plot of generated and training data for MiDi, EDM, and FMG for the GEOM and QM9 datasets. All models capture these conformational properties very well on QM9, and reasonably well on GEOM.
@@ A.2 Conditional generation
We train classifiers $\phi_{c}$ for the two considered properties, $\alpha$ and $\Delta\epsilon$. The property classifiers $\phi_{c}$ are E(3) invariant graph neural networks that receive pairwise distances between atoms $||\mathbf{m}_{i}-\mathbf{m}_{j}||$ and atom types as categorical features, very similar to the architecture used for parametrizing EDM. Please refer to (Guan et al., 2023) for more details.
We test our generative models by predicting the generated molecules’ properties, and computing the L1 loss between those predictions and the variable we used to condition the method during generation, $c$.
[…]
We also include additional baselines: FMGuc is an unconditional model, where we sample molecules according to an appropriate number of atoms $p(N|c)$, and FMGrnd are completely random samples, from an unconditional model.

### v2

@@ 6.5 Ablation
Model Val(%) Atom(%) Mol(%) TVa(↓) TVb(↓)
@@ 7 Conclusion
We demonstrated that continuous field-based molecular representations can achieve highly competitive results. Through theoretical and empirical analysis, we confirmed that E(3) equivariant methods consistently generate enantiomers with equal probabilities. This is crucial, as incorrect specification of enantiomers can lead to a loss of drug effectiveness or severe side effects. We also discuss in detail why developing a corrected SE(3) invariant point-cloud method would be significantly more challenging and show that our alternative successfully generates correct chiral properties.
Limitations and future work.
While point-cloud methods scale quadratically with the number of atoms, our method depends directly on the size and shape of the molecules. For any new dataset, the dimensions of the field tensors will always dictate the method’s computational feasibility and should be carefully considered.
Our tensor dimensions remain fairly large (even with our coarse resolution). Adaptations of current solutions of improved scaling (Gu et al., 2024; Ramesh et al., 2022; Rombach et al., 2022) could, in principle, allow for much larger architectures. In turn, scaling at the expense of group invariances may result in significant strides towards drug-like molecule design (Abramson et al., 2024).
We show significant performance gains using reference rotations, keeping general graph metrics competitive to E(3) formulations while obtaining chiral sensitivity. It remains, however, an open question whether our method can be applied to other 3D object generation tasks.
@@ 8 Acknowledgements
We acknowledge the computational resources provided by the Aalto Science-IT Project and CSC–IT Center for Science, Finland. This work was supported by Research Council of Finland (grants 334600, 359135, 342077), the Jane and Aatos Erkko Foundation (grant 7001703), and the Cancer Foundation of Finland.
@@ A.1 Conformational analysis
We compute the Wasserstein W1 distances for bond angle (BA), and bond lengths (BL) and respectively averaged over atom and bond types. We additionally compute conformational energies (CEMMFF and CExTB) W1 (definitions of these metrics can be found in Appndix H). We report the W1 distances in Tables 4 and 5. In Figures 4 and 5, we also plot the CDF of the distributions, based on which the W1 distances were computed and in Figure 6 and 7 we report the histograms of xTB energies.
Note that for the GEOM-Drugs dataset, MiDi and JODO use explicitly generated aromatic bond types, and thus has a different bond type set $\mathcal{B}$ compared to ours. This also affects the distributions of the bond types 1 and 2 and we therefore omit MiDi from the GEOM-Drugs bond distance metric BL.
Fields cannot enforce the number of atoms $N$ that will be generated, while point clouds do so, sampling $N\sim p_{\mathcal{D}}(N)$ during generation. While not enforcing the number of atoms provides useful flexibility (e.g., in conditional generation, the conditioning may be $p(\cdot|y)$, instead of $p(\cdot|y,N)$, where the latter requires knowledge of the joint $p(y,N)$), we find TVa (ablation Table 3) and CExTB (Table 4) to gain significant improvements as a result of atom count conditioning (from no conditioning, FMGβ=0, to FMGβ=2 which was used in this paper), but in this work, we opted for classifier-free guidance on binned atom intervals (see Appendix G), rather than conditioning on a specific atom count. The high impact of atom count conditioning on these metrics provides a strong indication that, if more fine-grained conditioning is adopted, fields will achieve similar TVa and CExTB numbers as point-clouds.
Table 4: Conformation results of molecules generated by models trained on the QM9 dataset, based on 10k samples. We report the mean and standard deviation of the metrics for three different runs (each generating 10k samples). The remark column (R) symbols represent: (*) re-tested by us. Reported bond length W1 distances (BL) are multiplied by $100$.
Model R BL (↓W1) BA (↓W1) CEMMFF (↓W1) CExTB (↓W1)
EDM (*) 0.47 0.68 3.45 0.30
MiDi (*) 1.92 0.95 3.12 0.34
JODO (*) 0.62 0.83 1.41 0.22
FMG - 0.59$\pm$0.01 1.08 $\pm$ 0.12 2.15$\pm$ 0.385 0.71±0.02
FMGβ=0 - - - - 0.98
Table 5: Conformation results of molecules generated by models trained on the GEOM-Drugs, based on 10k samples. The remark column (R) symbols represent: (*) re-tested by us ; ($\dagger$) post-processing corrections or algorithmic bond extraction; ()($\ddagger$)( ‡ ) explicit aromatic bonds. Reported bond length W1 distances (BL) are multiplied by $10^{2}$.
Model R BL (↓W1) BA (↓W1) CEMMFF (↓W1) CExTB (↓W1)
EDM (*) 8.84 4.36 - 1.93
MiDi (*$\ddagger$) - 1.38 40.98 1.82
JODO (*$\ddagger$) - 0.351 7.59 1.21
FMG - 2.74 1.00 53.14 3.22
Figure 4: Cumulative distribution function plot of generated and training data for MiDi, EDM, and FMG for the GEOM-Drugs and QM9 datasets. All models capture these conformational properties very well on QM9, and reasonably well on GEOM.
Figure 5: Cumulative distribution function plot of generated and training data energy conformations (kcal/mol) for MiDi, EDM, and FMG for the GEOM-Drugs and QM9 datasets. On GEOM, not enough valid samples were obtained to compute the energy distribution. Both models generate higher energies than the data on GEOM-Drugs (notably, MiDi was trained on a subset of lower energy molecules on GEOM).
Figure 6: xTB energy conformations (Eh) for MiDi, EDM, JODO, and FMG for the QM9 dataset.
Figure 7: xTB energy conformations (Eh) for MiDi, EDM, JODO, and FMG for the GEOM-Drugs dataset.
@@ A.2 Conditional generation
We train classifiers $\phi_{c}$ for the two considered properties, $\alpha$ and $\Delta\epsilon$. The property classifiers $\phi_{c}$ are E(3) invariant graph neural networks that receive pairwise distances between atoms $||\mathbf{m}_{i}-\mathbf{m}_{j}||$ and atom types as categorical features, similar to the architecture used for parametrizing EDM. We follow the setting from (Hoogeboom et al., 2022), splitting the 100 thousand molecules into two equal subsets, and training our generative method and property classifiers on each of the two subsets, respectively.
To sample the properties $c\sim p_{\mathcal{D}}$, we split molecular properties in one thousand equal bins, and form, based on training data statistics, the conditional and marginal categorical distributions $p(N|\mathbf{c}_{i})$, and $p(\mathbf{c}_{i})$, where $\mathbf{c}_{i}$ is the $i^{\text{th}}$ property bin and $N$ is the number of atoms. We test our conditional model by sampling two thousand pairs of $[\mathbf{c}_{i},N]\sim p(\mathbf{c}_{i})p(N|\mathbf{c}_{i})$, and further sample the actual conditional property $c\sim\mathcal{U}(\mathbf{c}_{i,0},\mathbf{c}_{i,1})$, with which we condition the generative model on $\bm{\epsilon}_{\theta}(\mathbf{f}_{t},t,N_{a},c)$.
We test our generative models by predicting the generated molecules’ properties, and computing the L1 loss between those predictions and the conditioning variable $c$ used during generation.
In Table 6, FMG is the trained conditional method for the two properties, respectively, and FMGβ=0 is the same model, where we do not use classification-free guidances on the number of atoms according to the property $p(N|c)$.
We also include additional baselines: FMGuc is an unconditional model, where we sample molecules according to an appropriate number of atoms $p(N_{a}|c)$, and FMGrnd are completely random samples, from an unconditional model.